# Satellite Autonomous Clock Fault Monitoring with Inter-Satellite Ranges Using Euclidean Distance Matrices

### v2

@@ Abstract
[Abstract] To address the need for robust positioning, navigation, and timing services in lunar environments, this paper proposes a novel onboard clock phase jump detection framework for satellite constellations using range measurements obtained from dual one-way inter-satellite links. Our approach leverages vertex redundantly rigid graphs to detect faults without relying on prior knowledge of satellite positions or clock biases, providing flexibility for lunar satellite networks with diverse satellite types and operators. We model satellite constellations as graphs, where satellites are vertices and inter-satellite links are edges. The proposed algorithm detects and identifies satellites with clock jumps by monitoring the singular values of the geometric-centered Euclidean distance matrix (GCEDM) of 5-clique sub-graphs. The proposed method is validated through simulations of a GPS constellation and a notional constellation around the Moon, demonstrating its effectiveness in various configurations.
@@ 1 Introduction
To meet the growing need for robust positioning, navigation, and timing (PNT) services at the lunar surface and lunar orbits, NASA and its international partners are collaborating to develop LunaNet (Israel2020), a network of networks providing data relay, PNT, detection, and science services. In LunaNet, these services are provided through cooperation among interoperable systems that evolve over time to meet the growing needs for these services, efficiently establishing a reliable, sustainable, and scalable network.
It is crucial to monitor LunaNet navigation satellites for the reliable operation of safety-critical missions, such as lunar landing and human missions on the lunar surface. The quality of lunar navigation signals can be compromised by various system faults, such as clock runoffs (e.g., phase and frequency jumps (weiss2010board)), unflagged maneuvers, failures in satellite payload signal generation components, and code-carrier incoherence (Walter2018). While the LunaNet Relay Service Documents (SRD) state that robustness of the navigation signal should be a key consideration for LunaNet (nasa2022srd), The specific methodology to monitor faults on LunaNet satellites has not yet been solidified.
The currently implemented integrity monitoring method for terrestrial GNSS can be categorized into three methods: (1) ground-based monitoring, (2) receiver autonomous integrity monitoring (RAIM), and (3) satellite-based integrity monitoring (SAIM). The first method relies on a network of ground-based monitoring stations at known positions to track and collect the GNSS signals, which are then collected and processed at a central computing center to compute corrections and integrity information (MISRA_1993). The system fault alerts are provided as integrity information from GNSS augmentation systems, such as the satellite-based augmentation system (SBAS) (van2009gps). The second method, RAIM, uses redundant measurements from GNSS satellites to detect and exclude faulty satellites on the receiver side (MISRA_1993). The third method, SAIM, computes the integrity information onboard the navigation satellite to detect satellite-related faults such as clock anomalies and ephemeris errors (xu2011gnss).
@@
The third method, SAIM, has several advantages compared with the other two methods: (1) faster time to alert enabled by eliminating the latency associated with data collection, processing, and uplink on the ground (2) the use of precise inter-satellite range (ISR) measurements (BeiDou satellites have demonstrated 10 cm accuracy (Zhang2024)) that are unaffected by the ionosphere and troposphere, and (3) the monitoring can be operated autonomously within the constellation and independently of ground monitoring. The autonomous nature of SAIM is particularly valuable for early lunar deployments, when lunar surface stations will be scarce or absent.
SAIM has been implemented in the BeiDou GNSS system to detect clock phase and frequency jumps by comparing the outputs of multiple onboard clocks (Cao2019; Chen2022), at the cost of adding extra hardware. Several other works have also proposed SAIM methods utilizing ISR measurements, which detect ephemeris anomalies by directly comparing the ISR measurement and the predicted range from the ephemeris  (wolf_onboard_2000; rodriguez-perez_inter-satellite_2011; xu2011gnss).
However, existing SAIM schemes process ISR measurements on a per-satellite basis, making their minimal detectable bias (MDB) dependent on ephemeris accuracy (xu2011gnss). This dependence may prove challenging for lunar constellations, where orbit determination and time synchronization are more difficult than for terrestrial GNSS. One remedy is to aggregate all ISR measurements centrally, form a linear system of residual equations, and apply data-snooping techniques (baarda1968testing; Imparato2018) to isolate faulty satellites. Yet this approach still requires the satellite positions to compute the residuals, which may not be (easily) available onboard all LunaNet nodes, since the interoperable LunaNet network may include links to satellites dedicated to communication or science tasks rather than PNT (Giordano2023).
In this paper, we propose a new algorithm to detect phase jumps in the satellite clocks that uses dual one-way ISR measurements. The proposed method operates directly on the observed ISRs and not the residuals, and thus does not require prior information or computation of the position or clock biases of the satellites. It detects faults by identifying biases injected into ISR measurements whenever a satellite’s onboard clock experiences a phase jump between two dual one-way links. We prove that we are able to detect fault satellites (satellites with clocks that experienced phase jumps) by assessing the realizability of the 5-clique (a fully connected graph of 5 nodes) subgraphs, since these subgraphs remain rigid (graphs that cannot be susceptible to continuous flexing are called rigid graphs), after removing any vertex from the graph (alireza_motevallian_robustness_2015). The concept of the algorithm is illustrated in Figure 1.
@@
Our proposed method monitors the singular values of the geometric-centered Euclidean distance matrix (GCEDM) (dokmanic2015euclidean) constructed from the ISR measurements, to check if 5-clique subgraphs are realizable in 3-dimensional space. In particular, we provide analytical expressions of the distribution of the squared 4th singular value under the presence of noise and biases in the range measurement, and design hypothesis tests to detect and identify the satellite clock with phase jumps.
The contribution of this paper is summarized below. This work is based on our prior conference paper presented at the 2024 ION GNSS+ Conference (Iiyama2024Rigid), but includes key theoretical and simulation updates beyond our prior publication.
  • We establish the necessary graph topology—specifically, vertex redundant rigidity—for unambiguous fault detection from ISR measurements.
@@
Note that our method is not intended to completely replace decades of research in the existing fault detection methods. Rather, our method provides new theoretical and practical insights to view the fault detection problem from a different perspective.
@@ 2.1 Dual One-Way Inter-satellite Range Measurement
We consider two satellites establishing a dual one-way inter-satellite link (ISL) as shown in Figure 2 to obtain ISR measurements. In this section, we formulate the ISR measurement model for the dual one-way link, based on Zhang2024.
@@
The corrections $\Delta\rho_{ij}$ and $\Delta\rho_{ji}$ are obtained using the predicted satellite ephemeris. Therefore, the accuracy of the corrections depends on the accuracy of the satellite velocity and the satellite clock drift of the ephemeris. For Beidou satellites, the accuracy of the predicted satellite velocity and clock drift is better than 0.1 mm/s and 1 ×$\times 10^{-13}$ s/s, resulting in $\Delta\rho_{ij},\Delta\rho_{ij}<1.0$ cm when $t_{2}-t_{1}<3$(Xie2022). For LunaNet satellites, the specification for the sum of velocity and clock drift error (converted to velocity by multiplying with lightspeed) in the ephemeris is 1.2 mm/s (for 3 $\sigma$), so the error of the correction terms can be in order of centimeters (nasa2022srd). Whenever further accuracy is required or velocity ephemeris information is not available, we can also utilize the inter-satellite range rate measurements between the satellites for correction (Iiyama2024contact).
@@ 2.2 Satellite Clock Phase Jumps
The (estimated) correction term will be
@@
$f_{i}$ =$=\begin{cases}\frac{c}{2}\cdot\delta t_{i}&\text{clock jump $\delta t_{i}$ occurs at satellite $i$ for $t$ within $t_{1}-\Delta t<t<t_{2}$}\\ 0&\text{satellite $i$'s clock is normal}\end{cases}$ (20)
$f_{j}$ =$=\begin{cases}\frac{c}{2}\cdot\delta t_{j}&\text{clock jump $\delta t_{j}$ occurs at satellite $j$ for $t$ within $t_{1}<t<t_{2}-\Delta t$}\\ 0&\text{satellite $j$'s clock is normal}\end{cases}$ (21)
@@ 3 Baseline Methods: Clock Fault Detection Using Inter-satellite Range Residuals
In this section, we introduce two baseline methods for detecting clock jump faults. The two methods uses the range residuals, which are computed from the ISL range measurements and the ephemeris information.
3.1 Fault Detection Using the Sum of Residuals
The first method is based on the ephemeris fault detection method proposed by rodriguez-perez_inter-satellite_2011 and xu2011gnss. While the original method was proposed for ephemeris fault detection, it could also be used for clock phase jump detection.
@@ 3.1 Fault Detection Using the Sum of Residuals
$T_{i}=\sum_{j=1}^{l_{i}}\left(\frac{g_{i,j}^{2}}{\sqrt{2\sigma_{r}^{2}+\sigma_{m}^{2}}}\right)^{2}$ (23)
@@
$\chi_{\alpha}^{2}(1,0)=P\left[\chi^{2}(1,0)\leq\chi_{\alpha}^{2}(1,0)\right]=\alpha$ (24)
where $\chi^{2}(r,\lambda)$ is the non-central Chi-squared distribution with $r$ degrees of freedom and the non-centrality parameter $\lambda$. The testing procedure is (xu2011gnss)
$\text{Accept}\ \mathcal{H}_{0}\ \text{if}\ \text{max}_{j\in\{1,\ldots,n\}}\ T_{i}<\sqrt{\chi_{\alpha}^{2}(1,0)}$ (25)
@@
The minimum detectable bias for $\mathcal{H}_{k}$ is computed as
@@
where the non-centrality parameter $\lambda(\alpha,\gamma,1)$ is computed as a function of false alarm rate $\alpha$ and detection power $\gamma$ (equal to the miss-detection rate of $1-\gamma$), by solving for the following equation
@@
Due to the correlation between $g_{i,j}$ and $g_{i,k}(i\neq k)$ via the common ephemeris error, the distribution of the test statistic under both hypotheses will not exactly follow a Chi-squared and non-central Chi-squared distribution, respectively. Based on this observation, xu2011gnss proposed to compute a conservative threshold for the test statistic via numerical simulation by using the correlation between two links $g_{i,j}$ and $g_{i,k}$ as follows
$\rho(g_{i,j},g_{i,k})\leq\frac{\sigma_{r}^{2}}{2\sigma_{r}^{2}+\sigma_{m}^{2}}$ (29)
In this paper, we used Imhof’s method (imhof1961computing) to compute the critical value for the test statistic. The detailed procedure is explained in Appendix A.
The advantage of this “sum of residuals” approach is that the fault detection is fully distributed among the satellites. The test is also able to detect the faults in the ephemeris information if they are present. However, the disadvantage of the fault detection method is that the minimum detectable bias is dependent on the ephemeris error $\sigma_{r}$, as shown in Equation (27).
@@ 3.2 Data Snooping with Baarda’s w-test Statistic
${c}_{f}^{(i)}=\begin{cases}1&\text{(satellite $f$ is transmitter of link $i$) \& (clock phase jump at satellite $f$ occurred during link $i$)}\\ -1&\text{(satellite $f$ is receiver of link $i$) \& (clock phase jump at satellite $f$ occurred during link $i$)}\\ 0&\text{otherwise}\end{cases}$ (31)
@@
For the test, we consider $n+1$ hypotheses: The null hypothesis $\mathcal{H}_{0}$ where the model errors are absent, and the alternative hypotheses $\mathcal{H}_{1},\ldots,\mathcal{H}_{n}$ which corresponds to the fault satellite being $f=1,f=2,\ldots,f=n$, respectively. The Baarda’s $w$-test statistic for testing $\mathcal{H}_{0}$ against single alternative $\mathcal{H}_{k}(k=1,\ldots,n)$ is given as (deJong2014)
$w_{k}=\frac{\mathbf{c}_{k}^{\top}\Sigma_{yy}^{-1}\mathbf{P_{A}}^{\perp}y}{\sqrt{\mathbf{c}_{k}^{\top}\Sigma_{yy}^{-1}\mathbf{P_{A}}^{\perp}\mathbf{c}_{k}}}$ (33)
where the projector $\mathbf{P_{A}}^{\perp}$ is defined as
$\mathbf{P_{A}}^{\perp}=\mathbf{I}_{m}-\mathbf{H}(\mathbf{H}^{\top}\mathbf{\Sigma_{yy}}^{-1}\mathbf{H})^{-1}\mathbf{H}^{\top}\mathbf{\Sigma_{yy}}^{-1}$ (34)
@@
${c}_{k}^{(i)}=\begin{cases}1&\text{satellite $k$ is transmitter of link $i$}\\ -1&\text{satellite $k$ is receiver of link $i$}\\ 0&\text{otherwise}\end{cases}$ (35)
Here, we assume that the clock phase jump at satellite $k$ corrupted all the ISLs that involve satellite $k$. The testing procedure is summarized as follows (Imparato2018).
@@
The minimum detectable bias for $\mathcal{H}_{k}$ is computed as (deJong2014)
$MDB(k)=\sqrt{\frac{\lambda(\alpha,\gamma,1)}{\mathbf{c}_{k}^{\top}\Sigma_{yy}^{-1}\mathbf{P_{A}}^{\perp}\mathbf{c}_{k}}}$ (38)
Iterating over $n$ alternate hypotheses using Baarda’s $w$-test statistic (which is called data snooping) enables the detection of the existence of biases by considering the effect of the biases on the full set of observations. Each test is optimal for any bias magnitude $b$, whenever the assumptions of the test are true (Lehmann2017). At the same time, it also have several limitations:
@@
    In $c_{k}$, we assume that the biases are injected into all ISR measurements that contains a satellite with a clock phase jump. However, this assumption is not always true since the clock jumps may have occurred outside the transmission and reception time for some links (the second condition in Equation (31) is not satisfied), based on the satellite link schedule (Iiyama2024contact). This will result in having smaller fault links than the assumed model ($\mathbf{c}_{k}$), and will result in the decreased detection power. If we want to rigorously handle this uncertainty of $\mathbf{c}_{k}$, we would need to consider all possible combinations of the links as alternate hypothesis, but this would be computationally intractable.
@@
    The (approximate) prior information of the satellite positions is needed to compute the Jacobian matrix $\mathbf{H}$. For navigation satellite constellations this is not often a problem since at least a coarse ephemeris is available. However, whenever this information is not available, we need to first perform the estimation of the satellite positions from the range measurements. This challenge is especially relevant for LunaNet: its interoperable design will use links to satellites serving other roles (e.g., data relay or scientific missions) that may not broadcast any ephemeris information.

### v3

@@ Abstract
[Abstract] To support reliable positioning, navigation, and timing (PNT) services for lunar satellite constellations, this paper proposes a framework for detecting satellite clock phase jumps using dual one-way inter-satellite range (ISR) measurements. The method models the constellation as a weighted graph and exploits rigidity properties to detect faults. Specifically, satellite clock faults are identified by monitoring the realizability of vertex-redundantly rigid subgraphs through the singular values of geometric-centered Euclidean distance matrices (GCEDMs). We establish the graph-topological conditions required for fault detection and derive the statistical distribution of the GCEDM test statistic under both nominal and faulty conditions. To address sparse link connectivity, we also introduce an ephemeris-augmented formulation that combines measured and estimated ranges and performs fault detection on fault-detectable subgraphs. Simulation results using a terrestrial GPS constellation and a notional lunar constellation demonstrate that the proposed approach provides robust fault detection under both dense and sparse measurement conditions.
keywords:
@@ 1 Introduction
To meet the growing need for robust positioning, navigation, and timing (PNT) services at the lunar surface and lunar orbits, NASA and its international partners are collaborating to develop LunaNet [14], a network of networks providing data relay, PNT, detection, and science services. In LunaNet, these services are provided through cooperation among interoperable systems that evolve over time to meet the growing needs for these services, efficiently establishing a reliable, sustainable, and scalable network.
It is crucial to monitor LunaNet navigation satellites for the reliable operation of safety-critical missions, such as lunar landing and human missions on the lunar surface. The quality of lunar navigation signals can be compromised by various system faults, such as clock runoffs (e.g., phase and frequency jumps [29]), unflagged maneuvers, failures in satellite payload signal generation components, and code-carrier incoherence [28]. While the LunaNet Relay Service Documents (SRD) state that robustness of the navigation signal should be a key consideration for LunaNet [24], the specific methodology to monitor faults on LunaNet satellites has not yet been solidified.
The currently implemented integrity monitoring method for terrestrial GNSS can be categorized into three methods: (1) ground-based monitoring, (2) receiver autonomous integrity monitoring (RAIM), and (3) satellite-based integrity monitoring (SAIM). The first method relies on a network of ground-based monitoring stations at known positions to track and collect the GNSS signals, which are then collected and processed at a central computing center to compute corrections and integrity information [23]. The system fault alerts are provided as integrity information from GNSS augmentation systems, such as the satellite-based augmentation system (SBAS) [27]. The second method, RAIM, uses redundant measurements from GNSS satellites to detect and exclude faulty satellites on the receiver side [23]. The third method, SAIM, computes the integrity information onboard the navigation satellite to detect satellite-related faults such as clock anomalies and ephemeris errors [32].
@@
The third method, SAIM, has several advantages compared with the other two methods: (1) faster time to alert enabled by eliminating the latency associated with data collection, processing, and uplink on the ground (2) the use of precise inter-satellite range (ISR) measurements (BeiDou satellites have demonstrated 10 cm accuracy [33]) that are unaffected by the ionosphere and troposphere, and (3) the monitoring can be operated autonomously within the constellation and independently of ground monitoring. The autonomous nature of SAIM is particularly valuable for early lunar deployments, when lunar surface stations will be scarce or absent.
SAIM has been implemented in the BeiDou GNSS system to detect clock phase and frequency jumps by comparing the outputs of multiple onboard clocks [3, 4], at the cost of adding extra hardware. Several other works have also proposed SAIM methods utilizing ISR measurements, which detect ephemeris anomalies by directly comparing the ISR measurement and the predicted range from the ephemeris [30, 25, 32]. However, these methods processes ISR measurements on a per-satellite basis, making their minimal detectable bias (MDB) dependent on ephemeris accuracy [32]. This dependence may prove challenging for lunar constellations, where orbit determination and time synchronization are more difficult than for terrestrial GNSS. Another potential SAIM algorithm that can solve this problem is to aggregate all ISR measurements centrally, form a linear system of residual equations, and apply data-snooping techniques [2, 13] to isolate faulty satellites.
In this paper, we propose a new algorithm to detect phase jumps in the satellite clocks that uses dual one-way ISR measurements. The proposed method detects faults by identifying biases injected into ISR measurements whenever a satellite’s onboard clock experiences a phase jump between two dual one-way links. We prove that we are able to detect fault satellites (satellites with clocks that experienced phase jumps) by assessing the realizability of the 5-clique (a fully connected graph of 5 nodes) subgraphs, since these subgraphs remain rigid (graphs that cannot be susceptible to continuous flexing are called rigid graphs), after removing any vertex from the graph [1]. The concept of the algorithm is illustrated in Figure 1.
@@
Our proposed method monitors the singular values of the geometric-centered Euclidean distance matrix (GCEDM) [7] constructed from the ISR measurements, to check if 5-clique subgraphs are realizable in 3-dimensional space. In particular, we provide analytical expressions of the distribution of the squared 4th singular value under the presence of noise and biases in the range measurement, and design hypothesis tests to detect and identify the satellite clock with phase jumps.
The contribution of this paper is summarized below. This work is based on our prior conference paper presented at the 2024 ION GNSS+ Conference [10], but includes key theoretical and simulation updates beyond our prior publication.
  • We establish the necessary graph topology, specifically, vertex redundant rigidity, for unambiguous fault detection from ISR measurements.
@@
Our method is not intended to replace decades of research in the existing fault detection methods. Rather, our method provides new theoretical and practical insights to view the fault detection problem from a different perspective.
@@ 2.1 Dual One-Way Inter-satellite Range Measurement
We consider two satellites establishing a dual one-way inter-satellite link (ISL) as shown in Figure 2 to obtain ISR measurements. In this section, we formulate the ISR measurement model for the dual one-way link, based on [33].
@@
The corrections $\Delta\rho_{ij}$ and $\Delta\rho_{ji}$ are obtained using the predicted satellite ephemeris. Therefore, the accuracy of the corrections depends on the accuracy of the satellite velocity and the satellite clock drift of the ephemeris. For Beidou satellites, the accuracy of the predicted satellite velocity and clock drift is better than 0.1 mm/s and 1 $\times 10^{-13}$ s/s, resulting in $\Delta\rho_{ij},\Delta\rho_{ij}<1.0$ cm when $t_{2}-t_{1}<3$ s [31]. For LunaNet satellites, the specification for the sum of velocity and clock drift error (converted to velocity by multiplying with lightspeed) in the ephemeris is 1.2 mm/s (for 3 $\sigma$), so the error of the correction terms can be in order of centimeters [24]. Whenever further accuracy is required or velocity ephemeris information is not available, we can also utilize the inter-satellite range rate measurements between the satellites for correction [11].
@@ 2.2 Satellite Clock Phase Jumps
The calculated correction term will be
@@
[…]
@@ 3 Baseline Methods: Clock Fault Detection Using Inter-satellite Range Residuals
In this section, we introduce two baseline methods for detecting clock jump faults based on existing literature. The two methods use the range residuals, which are computed from the ISL range measurements and the ephemeris information.
3.1 Fault Detection Using Expected Range from Ephemeris
The first method is based on the ephemeris fault detection method proposed by [25] and [32]. While the original method was proposed for ephemeris fault detection, it could also be used for clock phase jump detection.
@@ 3.1 Fault Detection Using Expected Range from Ephemeris
$T_{i}=\sum_{j=1}^{l_{i}}\left(\frac{g_{ij}^{2}}{\sqrt{2\sigma_{r}^{2}+\sigma_{m}^{2}}}\right)^{2}$ (23)
@@
$\chi_{\alpha}^{2}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}l_{i}},0)=P\left[\chi^{2}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}l_{i}},0)\leq\chi_{\alpha}^{2}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}l_{i}},0)\right]=\alpha$ (24)
where $\chi^{2}(r,\lambda)$ is the non-central Chi-squared distribution with $r$ degrees of freedom and the non-centrality parameter $\lambda$. Based on [32], the testing procedure is
$\text{Accept}\ \mathcal{H}_{0}\ \text{if}\ \text{max}_{j\in\{1,\ldots,n\}}\ T_{i}<\sqrt{\chi_{\alpha}^{2}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}l_{i}},0)}$ (25)
@@
The minimum detectable bias, $MDB(k)$, for $\mathcal{H}_{k}$ is computed as
@@
where the non-centrality parameter $\lambda(\alpha,\gamma,{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}l_{i}})$ is computed as a function of false alarm rate $\alpha$ and detection power $\gamma$ (equal to the miss-detection rate of $1-\gamma$), by solving for the following equation
@@
Due to the correlation between $g_{i,j}$ and $g_{i,k}(i\neq k)$ via the common ephemeris error, the distribution of the test statistic under both hypotheses will not exactly follow a Chi-squared and non-central Chi-squared distribution, respectively. Based on this observation, [32] proposed to compute a conservative threshold for the test statistic via numerical simulation by using the correlation between two links $g_{i,j}$ and $g_{i,k}$ as follows
$\rho(g_{ij},g_{ik})\leq\frac{\sigma_{r}^{2}}{2\sigma_{r}^{2}+\sigma_{m}^{2}}$ (29)
In this paper, we used Imhof’s method [12] to compute the critical value for the test statistic. The detailed procedure is explained in Appendix A.
The advantage of this “ephemeris comparison” approach is that the fault detection is fully distributed among the satellites. In addition, the method does not require a particular geometry of links to perform fault detection: if the ephemeris accuracy is sufficiently high, the fault can be detected by directly comparing the expected and measured range. The test is also able to detect the faults in the ephemeris information if they are present. However, the disadvantage of the fault detection method is that the minimum detectable bias is dependent on the ephemeris error $\sigma_{r}$, as shown in Equation (27).
@@ 3.2 Data Snooping with Baarda’s w-test Statistic
[…]
@@
For the test, we consider $n+1$ hypotheses: The null hypothesis $\mathcal{H}_{0}$ where the model errors are absent, and the alternative hypotheses $\mathcal{H}_{1},\ldots,\mathcal{H}_{n}$ which corresponds to the fault satellite being $f=1,f=2,\ldots,f=n$, respectively. The Baarda’s $w$-test statistic for testing $\mathcal{H}_{0}$ against single alternative $\mathcal{H}_{k}(k=1,\ldots,n)$ is given as [6]
$w_{k}=\frac{\mathbf{c}_{k}^{\top}\Sigma_{yy}^{-1}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{P}_{\mathbf{A}}^{\perp}}y}{\sqrt{\mathbf{c}_{k}^{\top}\Sigma_{yy}^{-1}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{P}_{\mathbf{A}}^{\perp}}\mathbf{c}_{k}}}$ (33)
where the projector $\mathbf{P}_{\mathbf{A}}^{\perp}$ is defined as
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{P}_{\mathbf{A}}^{\perp}}=\mathbf{I}_{m}-\mathbf{H}(\mathbf{H}^{\top}\mathbf{\Sigma_{yy}}^{-1}\mathbf{H})^{-1}\mathbf{H}^{\top}\mathbf{\Sigma_{yy}}^{-1}$ (34)
@@
[…]
Here, we assume that the clock phase jump at satellite $k$ corrupted all the ISLs that involve satellite $k$. The testing procedure is summarized as follows [13].
@@
The minimum detectable bias for $\mathcal{H}_{k}$ is computed as [6]
$MDB(k)=\sqrt{\frac{\lambda(\alpha,\gamma,1)}{\mathbf{c}_{k}^{\top}\Sigma_{yy}^{-1}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{P}_{\mathbf{A}}^{\perp}}\mathbf{c}_{k}}}$ (38)
Iterating over $n$ alternate hypotheses using Baarda’s $w$-test statistic (which is called data snooping) enables the detection of the existence of biases by considering the effect of the biases on the full set of observations. Each test is optimal for any bias magnitude $b$, whenever the assumptions of the test are true [21]. At the same time, it also has several limitations:
@@
    In $c_{k}$, we assume that the biases are injected into all ISR measurements that contain a satellite with a clock phase jump. However, this assumption is not always true since the clock jumps may have occurred outside the transmission and reception time for some links (the second condition in Equation (31) is not satisfied), based on the satellite link schedule [11]. This will result in having smaller fault links than the assumed model ($\mathbf{c}_{k}$), and will result in decreased detection power. If we want to rigorously handle this uncertainty of $\mathbf{c}_{k}$, we would need to consider all possible combinations of the links as alternate hypothesis, but this would be computationally intractable.
@@
    As further discussed in Section 4.1, the method fundamentally detects faults by checking the consistency among redundant measurements. Therefore, a sufficient level of measurement redundancy is required. When the measurement graph is sparse, and redundancy cannot be guaranteed, certain fault modes become unobservable, and the method may fail to detect them. In these cases, it may be necessary to rely on direct comparison with ephemeris-based ranges, which do not depend on internal measurement consistency.
Despite these limitations, there remains significant room for methodological improvement to mitigate them. For example, instead of formulating $n$ hypotheses at the satellite (vertex) level, one could construct hypotheses at the edge (range) level, explicitly modeling which individual ISRs are corrupted. A subsequent inference layer could then estimate the faulty satellites from the set of detected faulty edges. Such a hierarchical formulation would better accommodate partial-link corruption. We leave the development and evaluation of these extensions to future work.